# disposition

### v2

@@ 4 Nonlinear Cone Separation Conditions
  • $-\operatorname{cl}(\operatorname{conv}\,K)$ is representable as a Bishop-Phelps cone (cf. Example 2.4).
Thus, under assumption (4.3) (with $\psi(\cdot):=||\cdot||$), there is a pair $(x^{*},\alpha)\in Y^{*}\times\operatorname{\mathbb{P}}$ and an equivalent norm $||\cdot||^{\prime}\sim||\cdot||$ such that $-\operatorname{cl}(\operatorname{conv}\,K)=C_{||\cdot||^{\prime}}(x^{*},\alpha)$. By Lemma 4.4, in the real normed space $(Y,||\cdot||^{\prime})$ with dimension two or higher we have $\operatorname{cl}\,S_{-\operatorname{cl}(\operatorname{conv}\,K)}(||\cdot||^{\prime})=\{x\in\operatorname{\mathbb{B}}_{||\cdot||^{\prime}}\mid x^{*}(x)\geq\alpha\}$, hence (4.3) and (4.4) with $\psi(\cdot):=||\cdot||^{\prime}$ and $\operatorname{cl}(\operatorname{conv}\,K)$ in the role of $K$ are valid. Let us finish the proof by assuming a) and b), respectively:
a) In view of Lemma 4.9 ($1^{\circ}$), $B_{A}(||\cdot||^{\prime})$ is compact if $B_{A}(||\cdot||)$ is compact. Proposition 4.1 b) applied for the space $(Y,||\cdot||^{\prime})$, $\psi(\cdot):=||\cdot||^{\prime}$ and $\operatorname{cl}(\operatorname{conv}\,K)$ in the role of $K$ ensures $(\operatorname{cl}\,S_{A}^{0}(||\cdot||^{\prime}))\cap(\operatorname{cl}\,S_{-\operatorname{cl}(\operatorname{conv}\,K)}(||\cdot||^{\prime}))=\emptyset$, hence (4.1) with $\psi(\cdot):=||\cdot||^{\prime}$ is valid, taking into account that $\operatorname{cl}\,S_{{-\mathrm{c}l}(\operatorname{conv}\,K)}(||\cdot||^{\prime})\supseteq\operatorname{cl}\,S_{-K}(||\cdot||^{\prime})$.
b) In view of Lemma 4.9 ($2^{\circ}$, $4^{\circ}$), $\operatorname{cl}_{w}\,B_{A}(||\cdot||^{\prime})$ is weakly compact if either $Y$ is reflexive or $\operatorname{cl}_{w}\,B_{A}(||\cdot||)$ is weakly compact with $0\notin\operatorname{cl}_{w}\,B_{A}(||\cdot||)$. Proposition 4.1 b) applied for the space $Y$ endowed with the weak topology, $\psi(\cdot):=||\cdot||^{\prime}$ and $\operatorname{cl}(\operatorname{conv}\,K)$ in the role of $K$ ensures (4.1) with $\psi(\cdot):=||\cdot||^{\prime}$, taking into account that $\operatorname{cl}_{w}\,S_{A}^{0}(||\cdot||^{\prime})=\operatorname{cl}\,S_{A}^{0}(||\cdot||^{\prime})$ and $\operatorname{cl}_{w}\,S_{-\operatorname{cl}_{w}(\operatorname{conv}\,K)}(||\cdot||^{\prime})=\operatorname{cl}\,S_{{-\mathrm{c}l}(\operatorname{conv}\,K)}(||\cdot||^{\prime})\supseteq\operatorname{cl}\,S_{-K}(||\cdot||^{\prime})$.
@@
$A\cap-\operatorname{int}\,C_{\psi}(x^{*},\alpha)=\emptyset\quad\text{and}\quad K\subseteq C_{\psi}(x^{*},\alpha)\quad$ (4.11)
@@
  • $3^{\circ}$.
    There exists $(x^{*},\alpha)\in K^{a\circ}$ such that (4.11) is valid.
Proposition 4.15 ([28, Th. 5.9]).
Assume that $Y$ is a real locally convex space, $K,A\subseteq Y$ are nontrivial cones with normlike-bases $B_{K}$ and $B_{A}$. Suppose that one of the sets ${\mathrm{c}l}\,S_{-K}$ and ${\mathrm{c}l}\,S_{A}^{0}$ is compact, and (4.1) is valid. Then, there exists $(x^{*},\alpha)\in K^{a\#}\cap(Y^{*}\times\mathbb{P})$ such that
@@
$A\cap-C_{\psi}(x^{*},\alpha)=\{0\}\quad\text{and}\quad K\setminus\{0\}\subseteq C_{\psi}^{>}(x^{*},\alpha)=$ (4.12)
@@
From the general strict cone separation result in Proposition 4.15 and our Propositions 4.1 and 4.11 we immediately derive some new variants of strict cone separation results involving the classical separation condition (4.2).
Corollary 4.16.
@@
Assume that $Y$ is a real locally convex space, $K,A\subseteq Y$ are nontrivial cones with normlike-bases $B_{K}$ and $B_{A}$. Suppose that one of the sets ${\mathrm{c}l}\,S_{-K}$ and ${\mathrm{c}l}\,S_{A}^{0}$ is compact, and (4.2) and (4.3) are valid. Moreover, assume that the assumptions in a) or in b) of Proposition 4.1 are satisfied. Then, there is $(x^{*},\alpha)\in K^{a\#}\cap(Y^{*}\times\mathbb{P})$ such that (LABEL:sep_conditions_strict) is valid.
Corollary 4.17.

### v3

@@ 4.2 Nonlinear Cone Separation Conditions
  • $-\operatorname{cl}(\operatorname{conv}\,K)$ is representable as a Bishop-Phelps cone (cf. Example 1).
Thus, under assumption (12) (with $\psi(\cdot):=||\cdot||$), there is a pair $(x^{*},\alpha)\in Y^{*}\times\operatorname{\mathbb{P}}$ and an equivalent norm $||\cdot||^{\prime}\sim||\cdot||$ such that $-\operatorname{cl}(\operatorname{conv}\,K)=C_{||\cdot||^{\prime}}(x^{*},\alpha)$. By Lemma 20, in the real normed space $(Y,||\cdot||^{\prime})$ with dimension two or higher we have $\operatorname{cl}\,S_{-\operatorname{cl}(\operatorname{conv}\,K)}(||\cdot||^{\prime})=\{x\in\operatorname{\mathbb{B}}_{||\cdot||^{\prime}}\mid x^{*}(x)\geq\alpha\}$, hence (12) and (14) with $\psi(\cdot):=||\cdot||^{\prime}$ and $\operatorname{cl}(\operatorname{conv}\,K)$ in the role of $K$ are valid. Let us finish the proof by assuming a) and b), respectively:
a) In view of Lemma 22 ($1^{\circ}$), $B_{A}(||\cdot||^{\prime})$ is compact if $B_{A}(||\cdot||)$ is compact. Proposition 19 b) applied for the space $(Y,||\cdot||^{\prime})$, $\psi(\cdot):=||\cdot||^{\prime}$ and $\operatorname{cl}(\operatorname{conv}\,K)$ in the role of $K$ ensures $(\operatorname{cl}\,S_{A}^{0}(||\cdot||^{\prime}))\cap(\operatorname{cl}\,S_{-\operatorname{cl}(\operatorname{conv}\,K)}(||\cdot||^{\prime}))=\emptyset$, hence (10) with $\psi(\cdot):=||\cdot||^{\prime}$ is valid, taking into account that $\operatorname{cl}\,S_{{-\rm cl}(\operatorname{conv}\,K)}(||\cdot||^{\prime})\supseteq\operatorname{cl}\,S_{-K}(||\cdot||^{\prime})$.
b) In view of Lemma 22 ($2^{\circ}$, $4^{\circ}$), $\operatorname{cl}_{w}\,B_{A}(||\cdot||^{\prime})$ is weakly compact if either $Y$ is reflexive or $\operatorname{cl}_{w}\,B_{A}(||\cdot||)$ is weakly compact with $0\notin\operatorname{cl}_{w}\,B_{A}(||\cdot||)$. Proposition 19 b) applied for the space $Y$ endowed with the weak topology, $\psi(\cdot):=||\cdot||^{\prime}$ and $\operatorname{cl}(\operatorname{conv}\,K)$ in the role of $K$ ensures (10) with $\psi(\cdot):=||\cdot||^{\prime}$, taking into account that $\operatorname{cl}_{w}\,S_{A}^{0}(||\cdot||^{\prime})=\operatorname{cl}\,S_{A}^{0}(||\cdot||^{\prime})$ and $\operatorname{cl}_{w}\,S_{-\operatorname{cl}_{w}(\operatorname{conv}\,K)}(||\cdot||^{\prime})=\operatorname{cl}\,S_{{-\rm cl}(\operatorname{conv}\,K)}(||\cdot||^{\prime})\supseteq\operatorname{cl}\,S_{-K}(||\cdot||^{\prime})$. ∎
@@
$A\cap-\operatorname{int}\,C_{\psi}(x^{*},\alpha)=\emptyset\quad\text{and}\quad K\subseteq C_{\psi}(x^{*},\alpha)\quad$ (21)
@@
  • $3^{\circ}$
    There exists $(x^{*},\alpha)\in K^{a\circ}$ such that (21) is valid.
Proposition 26 ([14, Th. 5.9]).
Assume that $Y$ is a real locally convex space, $K,A\subseteq Y$ are nontrivial cones with normlike-bases $B_{K}$ and $B_{A}$. Suppose that one of the sets ${\rm cl}\,S_{-K}$ and ${\rm cl}\,S_{A}^{0}$ is compact, and (10) is valid. Then, there exists $(x^{*},\alpha)\in K^{a\#}\cap(Y^{*}\times\mathbb{P})$ such that
@@
$A\cap-C_{\psi}(x^{*},\alpha)=\{0\}\quad\text{and}\quad K\setminus\{0\}\subseteq C_{\psi}^{>}(x^{*},\alpha)=\operatorname{cor}\,C_{\psi}(x^{*},\alpha).$ (22)
@@
From the general strict cone separation result in Proposition 26 and our Propositions 19 and 23 we immediately derive some new variants of strict cone separation results involving the classical separation condition (11).
Corollary 27.
@@
Assume that $Y$ is a real locally convex space, $K,A\subseteq Y$ are nontrivial cones with normlike-bases $B_{K}$ and $B_{A}$. Suppose that one of the sets ${\rm cl}\,S_{-K}$ and ${\rm cl}\,S_{A}^{0}$ is compact, and (11) and (12) are valid. Moreover, assume that the assumptions in a) or in b) of Proposition 19 are satisfied. Then, there is $(x^{*},\alpha)\in K^{a\#}\cap(Y^{*}\times\mathbb{P})$ such that (22) is valid.
Corollary 28.